# The SKA Galactic Centre Survey A White Paper

### v1

@@ A White Paper
Rainer Schödel Instituto de Astrofísica de Andalucía (CSIC), Granada, Spain Antxón Alberdi Instituto de Astrofísica de Andalucía (CSIC), Granada, Spain Izaskun Jiménez-Serra Centro de Astrobiología (CAB), CSIC-INTA, Torrejón de Ardoz, Spain Farhad Yusef-Zadeh Northwestern University, Evanston, IL, USA Angela Gardini Instituto de Astrofísica de Andalucía (CSIC), Granada, Spain Michael Kramer Max-Planck-Institute for Radio Astronomy, Bonn, Germany Miguel Pérez-Torres Instituto de Astrofísica de Andalucía (CSIC), Granada, Spain Mark R. Morris University of California Los Angeles, Los Angeles, CA, USA Jan Forbrich University of Hertfordshire, Centre for Astrophysics Research, College Lane, Hatfield AL10 9AB, UK Adriano Ingallinera INAF - Osservatorio Astrofisico di Catania, Catania, Italy Francisco Nogueras-Lara European Southern Observatory (ESO), Garching bei München, Germany Jonathan D. Henshaw Liverpool John Moores University, Liverpool, UK Steven N. Longmore Liverpool John Moores University, Liverpool, UK Javier Moldón Instituto de Astrofísica de Andalucía (CSIC), Granada, Spain Ian Heywood University of Oxford, Oxford, UK Rhodes University, Makhanda, South Africa South African Radio Astronomy Observatory, South Africa Isabella Rammala Max-Planck-Institute for Radio Astronomy, Bonn, Germany Lourdes Verdes Montenegro Instituto de Astrofísica de Andalucía (CSIC), Granada, Spain Susana Sánchez Expósito Instituto de Astrofísica de Andalucía (CSIC), Granada, Spain
(June 6, 2024)
@@ 2 Science Goals
    What is the origin of the magnetized radio filaments in the GC?
@@ 5.6 Origin of the non-thermal filaments
One of the first hints showing that the nucleus of our Galaxy harbored energetic activity was the discovery of the archetype magnetized radio filaments in the GC 40 years ago (Yusef-Zadeh et al. 1984). Since then, radio observations have shown a population of linearly polarized synchrotron emission tracing nucleus-wide cosmic ray activity throughout the inner few hundred parsecs of the Galaxy (e.g. Heywood et al. 2019). Furthermore, H+${}_{3}^{+}$ absorption-line measurements of this region determined high cosmic-ray ionization rates indicating that relativistic particles permeate the CMZ at levels a thousand times that in the solar neighborhood (Oka et al. 2005; Indriolo & McCall 2012; Le Petit et al. 2016). These particles provide a significant source of pressure in the GC when compared to thermal gas pressure in the ISM of the GC. It is also becoming clear that there is another population of radio filaments detected in radio galaxies in poor clusters. These extragalactic filaments show remarkably similar underlying physics to those of the GC, in spite of vastly different scale lengths and environments (Yusef-Zadeh et al. 2023). The origin of this class of filaments is not understood either (Ramatsoku et al. 2020; Rudnick et al. 2022). However, the similarities provide an opportunity to investigate the physical processes in the ISM of the GC and compare them to the environment of the intracluster medium for the first time using detailed and sensitive SKA observations.
@@ 5.8 Extinction towards the GC
Interstellar extinction is extreme and highly variable as a function of the line-of-sight towards the GC (see e.g. Scoville et al. 2003; Nogueras-Lara et al. 2020b). It is one of the greatest obstacles in our interpretation of infrared observations of stars in the GC. For example, accurate and precise measurements of extinction are needed to correct observed stellar magnitudes and thus to be able to classify stars approximately via photometric measurements (e.g. Schoedel et al. 2023). Also, the conversion of observed fluxes to absolute fluxes depends sensitively on an accurate knowledge of extinction, implying an insecurity of a factor of up to two in the estimated stellar masses of young, massive stars (e.g. Clark et al. 2018). Combining measurements of radio hydrogen recombination lines with infrared HI emission lines is a highly reliable way of measuring absolute extinction as a function of wavelength towards the GC (see Scoville et al. 2003; Fritz et al. 2011).
Emission from ionised ISM is pervasive in the GC. SKA observations of radio recombination lines, which are unaffected by extinction, can be combined with infrared recombination lines from NIRCam/JWST (see Schoedel et al. 2023). Such data would allow us to determine the extinction curve with great accuracy across the GC and test whether and how it depends on the line of sight and on infrared wavelength (see, e.g., discussion in Nogueras-Lara et al. 2021). In addition, knowledge of extinction towards different ISM features will allow us to constrain their relative distances along the line of sight and therefore the 3D-distribution of the ISM at the GC.
@@ 6 Properties of point-like radio sources at the GC
Figure 4 provides an overview of point-like sources at the GC and their expected flux densities at different frequencies. Apart from their fluxes, the different types of point sources can be distinguished in various ways: Wolf Rayet and O-type main and post main sequence stars will have bright near-infrared counterparts and frequently thermal radio spectra. Pulsars will typically have steep spectra (with the possible exception of magnetars, which may have flat to inverted spectra, see Zhao et al. 2022), polarized emission (Sobey et al. 2022) and show little variability. LMXBs will show great variability with most of them manifesting as transient sources. In the low/hard state they typically show flat to inverted spectra from optically thick jet bases, i.e. $\alpha>0$ for $S_{\nu}\propto\lambda^{\alpha}$, where $S_{\nu}$ is the radio flux density and $\lambda$ the wavelength, but they become optically thin ($\alpha<0$) in the high/soft state (Fender et al. 2004). As concerns multi-wavelength cross-identification, Ultracompact HII and similar nebular sources have sizes $\lesssim 0.1$ pc, corresponding to $\lesssim 2.5"$ at the distance of the GC. Practically all of them will therefore be resolved and/or have an infrared counterpart. Near-infrared imaging with an angular resolution of $0.2"$ is available from the GALACTICNUCLEUS survey (Nogueras-Lara et al. 2019). Higher angular resolution imaging from JWST NIRCam has already been obtained on some selected fields (such as the central parsecs, Sgr C, or the Brick molecular cloud) and may become available for a significant fraction of the GC in the next years (Schoedel et al. 2023). Also, specific fields may be observed with adaptive optics assisted near-infrared instruments on large ground-based telescopes, such as ERIS/VLT.

### v2

@@ A White Paper
Rainer Schödel Instituto de Astrofísica de Andalucía (CSIC), Granada, Spain Antxón Alberdi Instituto de Astrofísica de Andalucía (CSIC), Granada, Spain Izaskun Jiménez-Serra Centro de Astrobiología (CAB), CSIC-INTA, Torrejón de Ardoz, Spain Farhad Yusef-Zadeh Northwestern University, Evanston, IL, USA Angela Gardini Instituto de Astrofísica de Andalucía (CSIC), Granada, Spain Michael Kramer Max-Planck-Institute for Radio Astronomy, Bonn, Germany Miguel Pérez-Torres Instituto de Astrofísica de Andalucía (CSIC), Granada, Spain Mark R. Morris University of California Los Angeles, Los Angeles, CA, USA Jan Forbrich University of Hertfordshire, Centre for Astrophysics Research, College Lane, Hatfield AL10 9AB, UK Adriano Ingallinera INAF - Osservatorio Astrofisico di Catania, Catania, Italy Francisco Nogueras-Lara European Southern Observatory (ESO), Garching bei München, Germany Jonathan D. Henshaw Liverpool John Moores University, Liverpool, UK Steven N. Longmore Liverpool John Moores University, Liverpool, UK Javier Moldón Instituto de Astrofísica de Andalucía (CSIC), Granada, Spain Ian Heywood University of Oxford, Oxford, UK Rhodes University, Makhanda, South Africa South African Radio Astronomy Observatory, South Africa Isabella Rammala Max-Planck-Institute for Radio Astronomy, Bonn, Germany Farideh Mazoochi Institute for Research in Fundamental Sciences (IPM), School of Astronomy, Tehran, Iran Fatemeh Tabatabaei Institute for Research in Fundamental Sciences (IPM), School of Astronomy, Tehran, Iran Lourdes Verdes Montenegro Instituto de Astrofísica de Andalucía (CSIC), Granada, Spain Susana Sánchez Expósito Instituto de Astrofísica de Andalucía (CSIC), Granada, Spain
(May 28, 2025)
@@ 2 Science Goals
    What are the properties and physics of the large-scale magnetic field in the GC?
@@ 5.5 Searching for the “building blocks” of life
5.6 Properties and physics of the large-scale magnetic field
Among the biggest open questions concerning the GC magnetic field are: (1) Is the milligauss vertical field homogeneous and pervasive throughout the CMZ and does this define an overall GC magnetosphere or are the strongest magnetic features only localised, transient features? (2) How and where does the vertical field couple to the horizontal field? Can we find specific points of interaction? (3) What is the origin of the relativistic electrons that light up the NTFs (see next subsection)? (4) What is the origin of the poloidal field?
Studies of MeerKAT continuum data at 1.3 GHz (e.g. Heywood et al. 2022; Yusef-Zadeh et al. 2023) have traced the thermal and non-thermal filaments with unprecedented sensitivity and have provided an impressive new view of the GC magnetic field. Still, they are limited to an angular resolution $\simeq 5^{"}$, which makes it hard to untangle different features and detect faint filaments in the radio bright GC.
The SKA1-Mid in bands 2, 5a, and 5b will provide sensitive, high-resolution radio continuum maps, enabling to study the magnetic field with angular resolutions almost ten times better than the published MeerKAT observations. This will shed light on the role of the magnetic field in the GC on scales down to milli-parsecs. So far undiscovered faint filaments will improve our knowledge of the magnetic field and of the origin of the relativistic electrons that power the NTFs. One particularly important question here is the existence or not of non-thermal horizontal filaments close to the Galactic plane. The data will provide us with a better understanding of the influence of the magnetic field on the dynamics of the ISM, which may even come to dominate some HII regions (see Bally:2024iu). An exciting prospect is that the high angular resolution of SKA-MID may allow us to perform proper motion measurements of magnetic filaments over timespans of a few to ten years. Proper motions could help us clarify the relative locations, interactions and whether the vertical field does or does not rotate with respect to the ISM and stars, which could indicate whether we are seeing global field configurations or whether the NTFs are localised, short-lived features.
@@ 5.7 Origin of the non-thermal filaments
One of the first hints showing that the nucleus of our Galaxy harboured energetic activity was the discovery of the archetype magnetised radio filaments in the GC 40 years ago (Yusef-Zadeh et al. 1984). Since then, radio observations have shown a population of linearly polarised synchrotron emission tracing nucleus-wide cosmic ray activity throughout the inner few hundred parsecs of the Galaxy (e.g. Heywood et al. 2019). Furthermore, H+${}_{3}^{+}$ absorption-line measurements of this region determined high cosmic-ray ionisation rates indicating that relativistic particles permeate the CMZ at levels a thousand times that in the solar neighbourhood (Oka et al. 2005; Indriolo & McCall 2012; Le Petit et al. 2016). These particles provide a significant source of pressure in the GC when compared to thermal gas pressure in the ISM of the GC. It is also becoming clear that there is another population of radio filaments detected in radio galaxies in poor clusters. These extragalactic filaments show remarkably similar underlying physics to those of the GC, in spite of vastly different scale lengths and environments (Yusef-Zadeh et al. 2023). The origin of this class of filaments is not understood either (Ramatsoku et al. 2020; Rudnick et al. 2022). However, the similarities provide an opportunity to investigate the physical processes in the ISM of the GC and compare them to the environment of the intracluster medium for the first time using detailed and sensitive SKA observations.
@@ 5.9 Extinction towards the GC
Interstellar extinction is extreme and highly variable as a function of the line-of-sight towards the GC (see e.g. Scoville et al. 2003; Nogueras-Lara et al. 2020b). It is one of the greatest obstacles in our interpretation of infrared observations of stars in the GC. For example, accurate and precise measurements of extinction are needed to correct observed stellar magnitudes and thus to be able to classify stars approximately via photometric measurements (e.g. Schoedel:2023rf). Also, the conversion of observed fluxes to absolute fluxes depends sensitively on an accurate knowledge of extinction, implying an insecurity of a factor of up to two in the estimated stellar masses of young, massive stars (e.g. Clark et al. 2018). Combining measurements of radio hydrogen recombination lines with infrared HI emission lines is a highly reliable way of measuring absolute extinction as a function of wavelength towards the GC (see Scoville et al. 2003; Fritz et al. 2011).
Emission from ionised ISM is pervasive in the GC. SKA observations of radio recombination lines, which are unaffected by extinction, can be combined with infrared recombination lines from NIRCam/JWST (see Schoedel:2023rf). Such data would allow us to determine the extinction curve with great accuracy across the GC and test whether and how it depends on the line of sight and on infrared wavelength (see, e.g., discussion in Nogueras-Lara et al. 2021). In addition, knowledge of extinction towards different ISM features will allow us to constrain their relative distances along the line of sight and therefore the 3D-distribution of the ISM at the GC.
@@ 6 Properties of point-like radio sources at the GC
Figure 4 provides an overview of point-like sources at the GC and their expected flux densities at different frequencies. Apart from their fluxes, the different types of point sources can be distinguished in various ways: Wolf Rayet and O-type main and post main sequence stars will have bright near-infrared counterparts and frequently thermal radio spectra. Pulsars will typically have steep spectra (with the possible exception of magnetars, which may have flat to inverted spectra, see Zhao et al. 2022), polarised emission (Sobey et al. 2022) and show little variability. LMXBs will show great variability with most of them manifesting as transient sources. In the low/hard state they typically show flat to inverted spectra from optically thick jet bases, i.e. $\alpha>0$ for $S_{\nu}\propto\lambda^{\alpha}$, where $S_{\nu}$ is the radio flux density and $\lambda$ the wavelength, but they become optically thin ($\alpha<0$) in the high/soft state (Fender et al. 2004). As concerns multi-wavelength cross-identification, Ultracompact HII and similar nebular sources have sizes $\lesssim 0.1$ pc, corresponding to $\lesssim 2.5"$ at the distance of the GC. Practically all of them will therefore be resolved and/or have an infrared counterpart. Near-infrared imaging with an angular resolution of $0.2"$ is available from the GALACTICNUCLEUS survey (Nogueras-Lara et al. 2019). Higher angular resolution imaging from JWST NIRCam has already been obtained on some selected fields (such as the central parsecs, Sgr C, or the Brick molecular cloud) and may become available for a significant fraction of the GC in the next years (Schoedel:2023rf). Also, specific fields may be observed with adaptive optics assisted near-infrared instruments on large ground-based telescopes, such as ERIS/VLT.